### v2

# Entropy-based random models for hypergraphs

Abstract
Network theory has primarily focused on pairwise relationships, disregarding many-body interactions: neglecting them, however, can lead to misleading representations of complex systems. Hypergraphs represent an increasingly popular alternative for describing polyadic interactions: our innovation lies in leveraging the representation of hypergraphs based on the incidence matrix for extending the entropy-based framework to higher-order structures. In analogy with the Exponential Random Graphs, we name the members of this novel class of models Exponential Random Hypergraphs. Here, we focus on two explicit examples, i.e. the generalisations of the Erdös-Rényi Model and of the Configuration Model. After discussing their asymptotic properties, we employ them to analyse real-world configurations: more specifically, i) we extend the definition of several network quantities to hypergraphs, ii) compute their expected value under each null model and iii) compare it with the empirical one, in order to detect deviations from random behaviours. Differently from currently available techniques, ours is analytically tractable, scalable and effective in singling out the structural patterns of real-world hypergraphs differing significantly from those emerging as a consequence of simpler, structural constraints.
pacs:
89.75.Fb; 02.50.Tt
I I. Introduction
Networks provide a powerful language to model interacting systems Newman (2010); Caldarelli (2010). Within a network framework, the basic unit of interaction, i.e. the edge, involves two nodes and the complexity of the structure as a whole arises from the combination of these units. Despite its many successes, network science disregards certain aspects of interacting systems, notably the possibility that more-than-two constituent units could interact at a time Lambiotte et al. (2019). Yet, it has been increasingly shown that, for a variety of systems, interactions cannot be always decomposed into a pairwise fashion and that neglecting higher-order ones can lead to an incomplete, if not misleading, representation of them Lambiotte et al. (2019); Battiston et al. (2020); Bick et al. (2023). Examples include chemical reactions involving several compounds, coordination activities within small teams of co-working people, brain activities mediated by groups of neurons. Generally speaking, thus, modelling the joint coordination of multiple entities calls for a framework that generalises the traditional, edge-centered one. An increasingly popular alternative to support a science of many-body interactions is provided by hypergraphs as these versatile, mathematical objects naturally allow nodes to interact in groups Battiston et al. (2021).
Given the recent interest towards these objects, the definition of analytical tools to study them is still in its infancy Chodrow (2020); Nakajima et al. (2021); Musciotto et al. (2021); Barthelemy (2022). This paper represents our contribution to fill the gap: hereby, we extend the class of entropy-based null models Jaynes (1957); Park and Newman (2004) to hypergraphs. These models work by preserving a given set of quantities while randomising everything else, hence destroying all, possible correlations between structural properties except for those that are genuinely embodied by the constraints themselves Squartini and Garlaschelli (2011); Cimini et al. (2018); Squartini et al. (2018). The versatility of such an approach allows it to be employed either in presence of full information (to quantify the level of self-organisation of a given configuration by spotting out the patterns that are incompatible with simpler, structural constraints Squartini et al. (2013); Saracco et al. (2017); Becatti et al. (2019); Caldarelli et al. (2020); Neal et al. (2021)) or in presence of partial information (to infer the missing portion of a given configuration Parisi et al. (2020)).
Our strategy for defining null models for hypergraphs is based on the randomisation of their incidence matrix. The (generally, rectangular) table contains information about the connectivity of nodes, i.e. the set of hyperedges they belong to, and the connectivity of hyperedges, i.e. the set of nodes they cluster. We will explicitly derive two members of the novel class of models hereby named Exponential Random Hypergraphs, i.e. the Random Hypergraph Model (RHM, generalising the Erdös-Rényi Model) and the Hypergraph Configuration Model (HCM, generalising the Configuration Model), and provide an analytical characterisation of their behaviour. To this aim, we will exploit the formal equivalence between the incidence matrix of a hypergraph and the biadjacency matrix of a bipartite graph Nakajima et al. (2021); Barthelemy (2022). As a next step, we will employ the HCM to assess the statistical significance of a number of patterns characterising several, real-word hypergraphs.
II II. Formalism and basic quantities
A hypergraph can be defined as a pair $H(\mathcal{V},\mathcal{E}_{H})$ where $\mathcal{V}$ is the set of vertices and $\mathcal{E}_{H}$ is the set of hyperedges. Moving from the observation that the edge set $\mathcal{E}_{G}$ of a traditional, binary, undirected graph $G(\mathcal{V},\mathcal{E}_{G})$ is a subset of the power set of $\mathcal{V}$, several definitions of the hyperedge set $\mathcal{E}_{H}$ have been provided: the two most popular ones are those proposed in Berge (1967, 1973), where hyperedges ‘link’ one or more vertices, and in Voloshin (2009), where hyperedges are allowed to be empty sets as well. Hereby, we adopt the definition according to which $\mathcal{E}_{H}$ is a multiset of the power set of $\mathcal{V}$: in words, we consider non-simple hypergraphs, admitting loops and parallel edges (i.e. hyperedges involving exactly the same set nodes) of any size, including 0 (corresponding to empty hyperedges), and $|\mathcal{V}|$ (corresponding to hyperedges clustering all vertices together).
$\mathbf{I}\>=\>\begin{tabular}[]{c|c|c|c|c|c|}&$\mathbf{e_{1}}$&$\mathbf{e_{2}}$&$\mathbf{e_{3}}$&$\mathbf{e_{4}}$&$\mathbf{e_{5}}$\\ \hline\cr$\mathbf{n_{1}}$&0&1&1&0&1\\ \hline\cr$\mathbf{n_{2}}$&0&0&1&1&1\\ \hline\cr$\mathbf{n_{3}}$&0&1&1&1&0\\ \hline\cr$\mathbf{n_{4}}$&1&0&0&1&1\\ \hline\cr$\mathbf{n_{5}}$&1&1&1&1&0\\ \hline\cr$\mathbf{n_{6}}$&0&1&0&0&0\\ \hline\cr\end{tabular}$
Table 1: Example of incidence matrix describing a binary, undirected hypergraph with $6$ nodes and $5$ hyperedges. The generic entry $I_{i\alpha}$ is 1 if vertex $i$ belongs to hyperedge $\alpha$ and 00 otherwise (see also Fig. 1).
As for traditional graphs, an algebraic representation of hypergraphs can be devised as well. For a formal analogy with the traditional case, let us call the cardinality of the set of nodes $|\mathcal{V}|=N$ and the cardinality of the set of hyperedges $|\mathcal{E}_{H}|=L$: then, we consider the $N\times L$ table known as incidence matrix, each row of which corresponds to a node and each column of which corresponds to a hyperedge. If we indicate the incidence matrix with $\mathbf{I}$, its generic entry $I_{i\alpha}$ will be $1$ if vertex $i$ belongs to hyperedge $\alpha$ and 00 otherwise. Notice that the number of $1$s along each row can vary between 00 and $L$, the former case indicating an isolated node and the latter one indicating a node that belongs to each hyperedge; similarly, the number of $1$s along each column can vary between 00 and $N$, the former case indicating an empty hyperedge and the latter one indicating a hyperedge that includes all nodes. As explicitly noticed elsewhere Nakajima et al. (2021); Barthelemy (2022), representing a hypergraph via its incidence matrix is equivalent at considering the bipartite graph defined by the sets $\mathcal{V}$ and $\mathcal{E}_{H}$. Table 1 represents the incidence matrix describing the binary, undirected hypergraph shown in Fig. 1.
Once the incidence matrix has been defined, several quantities needed for the description of hypergraphs can be defined quite straightforwardly: for example, the ‘degree of node $i$’ (hereby, degree) reads
$k_{i}=\sum_{\alpha=1}^{L}I_{i\alpha}$ (1)
and counts the number of hyperedges that are incident to it; analogously, the ‘degree of hyperedge $\alpha$’ (hereby, hyperdegree) reads
$h_{\alpha}=\sum_{i=1}^{N}I_{i\alpha}$ (2)
and counts the number of nodes it clusters. Both the sum of degrees and that of hyperdegrees equal the total number of $1$s, i.e. $\sum_{i=1}^{N}k_{i}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}=\sum_{\alpha=1}^{L}\sum_{i=1}^{N}I_{i\alpha}=\sum_{\alpha=1}^{L}h_{\alpha}\equiv T$. Importantly, a node degree no longer coincides with the number of its neighbours: instead, it matches the number of hyperedges it belongs to; a hyperdegree, instead, provides information on the hyperedge size.
Figure 1: Cartoon representation of the hypergraph whose incidence matrix is shown in Table 1.
III III. Binary, undirected hypergraphs randomisation
An early attempt to define randomisation models for hypergraphs can be found in Ghoshal et al. (2009). Its authors, however, have just considered hyperedges that are incident to triples of nodes - a framework that has been, later, applied to the study of the World Trade Network Nakajima et al. (2021). Considering the incidence matrix has two, clear advantages over the tensor-based representation employed in Ghoshal et al. (2009); Nakajima et al. (2021): i) generality, because the incidence matrix allows hyperedges of any size to be handled at once; ii) compactness, because the order of the tensor $\mathbf{I}$ never exceeds two, hence allowing any hypergraph to be represented as a traditional, bipartite graph.
In order to extend the rich set of null models induced by graph-specific, global and local constraints to hypergraphs, we, first, need to identify the quantities that can play this role within the novel setting. In what follows, we will consider the total number of 1s, i.e. $T$, the degree and the hyperdegree sequences, i.e. $\{k_{i}\}_{i=1}^{N}$ and $\{h_{\alpha}\}_{\alpha=1}^{L}$ either separately or in a joint fashion; moreover, we will distinguish between microcanonical and canonical randomisation techniques.
III.1 A. Homogeneous benchmarks:
the Random Hypergraph Model (RHM)
The model is defined by just one, global constraint, that, in our case, reads $T=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}$. Its microcanonical version extends the model by Erdös and Rényi Erdös and Rényi (1959) - also known as Random Graph Model (RGM) - to hypergraphs and prescribes to count the number of incidence matrices that are compatible with a given, total number of 1s, say $T^{*}$: they are
$\Omega_{\text{RHM}}=\binom{V}{T^{*}}$ (3)
with $V=NL$ being the total number of entries of the incidence matrix $\mathbf{I}$. Once the total number of configurations composing the microcanonical ensemble is determined, a procedure to generate them is needed: in the case of the RHM, it simply boils down to reshuffling the entries of the incidence matrix, a procedure ensuring that the total number of 1s is kept fixed while any other correlation is destroyed.
The canonical version of the RHM, instead, extends the model by Gilbert Gilbert (1959) and rests upon the constrained maximisation of Shannon entropy, i.e.
$\mathscr{L}\equiv S[P]-\sum_{i=0}^{M}\theta_{i}[P(\mathbf{I})C_{i}(\mathbf{I})-\langle C_{i}\rangle]$ (4)
where $S[P]=-\sum_{\mathbf{I}\in\mathcal{I}}P(\mathbf{I})\ln P(\mathbf{I})$, $C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining $M-1$ constraints represent proper, topological properties. The sum defining Shannon entropy runs over the set $\mathcal{I}$ of incidence matrices described in the introductory paragraph and known as canonical ensemble.
Such an optimisation procedure defines the Exponential Random Hypergraph (ERH) framework, described by the expression
$P(\mathbf{I})=\frac{e^{-H(\mathbf{I})}}{Z}=\frac{e^{-H(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-H(\mathbf{I})}}=\frac{e^{-\sum_{i=1}^{M}\theta_{i}C_{i}(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{M}\theta_{i}C_{i}(\mathbf{I})}}.$ (5)
In the simplest case, the only, global constraint is represented by $T$ and leads to the expression
$P(\mathbf{I})$ $=\frac{e^{-\theta T(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\theta T(\mathbf{I})}}$
$=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}$
$=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}x^{I_{i\alpha}}\prod_{i=1}^{N}\prod_{\alpha=1}^{L}(1+x)^{-1}$ (6)
that can be re-written as
$P(\mathbf{I})=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}p^{I_{i\alpha}}(1-p)^{1-I_{i\alpha}}=p^{T(\mathbf{I})}(1-p)^{NL-T(\mathbf{I})}$ (7)
with $e^{-\theta}\equiv x$ and $p\equiv x/(1+x)$. The canonical ensemble, now, includes all $N\times L$ rectangular matrices whose number of entries equating 1 ranges from 0 to $NL$. According to such a model, the entries of the incidence matrix are i.i.d. Bernoulli random variables, i.e. $I_{i\alpha}\sim\text{Ber}(p)$, $\forall\>i,\alpha$: as a consequence, the total number of 1s, the degrees and the hyperdegrees obey Binomial distributions - being all defined as sums of i.i.d. Bernoulli random variables. Specifically, $T\sim\text{Bin}(NL,p)$, $k_{i}\sim\text{Bin}(L,p)$, $\forall\>i$ and $h_{\alpha}\sim\text{Bin}(N,p)$, $\forall\>\alpha$, in turn, implying that $\langle T\rangle_{\text{RHM}}=NLp$, $\langle k_{i}\rangle_{\text{RHM}}=Lp$, $\forall\>i$ and $\langle h_{\alpha}\rangle_{\text{RHM}}=Np$, $\forall\>\alpha$.
In order to ensure that $\langle T\rangle_{\text{RHM}}=T^{*}$, parameters have to be tuned opportunely. To this aim, the likelihood maximisation principle can be invoked Garlaschelli and Loffredo (2008): it prescribes to maximise the function $\mathcal{L}(\theta)\equiv\ln P(\mathbf{I}^{*}|\theta)$ with respect to the unknown parameter(s) that define it. Such a recipe leads us to find $p=T^{*}/NL$, with $T^{*}=T(\mathbf{I}^{*})$ indicating the empirical value of the constraint defining the RHM.
The RHM (also considered in Barthelemy (2022), although a derivation from first principles is missing, there) is formally equivalent to the Bipartite Random Graph Model (BiRGM) Saracco et al. (2015). Such an identification is guaranteed by our focus on non-simple hypergraphs.
III.1.1 1. Estimation of the number of empty hyperedges
As non-simple hypergraphs admit the presence of empty as well as parallel hyperedges, let us evaluate how frequently they appear in the ensembles induced by our benchmarks.
To this aim, let us start by considering the probability for the generic hyperedge $\alpha$ to be empty or, in other terms, that its hyperdegree $h_{\alpha}$ is zero. Upon remembering that $h_{\alpha}\sim\text{Bin}(N,p)$, one finds
$p_{0}\equiv(1-p)^{N}.$ (8)
Such a quantity leads us to find the probability of observing at least one, empty hyperedge (or, in other terms, the complementary of the probability that no hyperedge is empty), i.e.
$p_{1}\equiv 1-(1-p_{0})^{L}=1-[1-(1-p)^{N}]^{L},$ (9)
and the expected number of empty hyperedges: upon remembering that they are i.i.d. Binomial random variables, one finds
$\langle N_{0}\rangle\equiv Lp_{0}=L(1-p)^{N}.$ (10)
Let us, now, inspect the behaviour of the aforementioned quantities on the ensemble induced by the RHM, as the density of 1s in the incidence matrix, i.e. $p=T/NL$, varies. The two regimes of interest are the dense one, defined by $T\rightarrow NL$, and the sparse one, defined by $T\rightarrow 0$. In the dense case, one finds
$\lim_{T\to NL}p_{0}$ $=\lim_{T\to NL}\left(1-\frac{T}{NL}\right)^{N}=0,$ (11)
a relationship inducing $p_{1}\xrightarrow[]{T\to NL}0$ and $\langle N_{0}\rangle\xrightarrow[]{T\to NL}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of 1s increases. Consistently, in the sparse case one finds
$\lim_{T\to 0}p_{0}=\lim_{T\to 0}\left(1-\frac{T}{NL}\right)^{N}=1,$ (12)
a relationship inducing $p_{1}\xrightarrow[]{T\to 0}1$ and $\langle N_{0}\rangle\xrightarrow[]{T\to 0}L$: in words, the probability of observing empty hyperedges progressively rises as the density of 1s decreases.
To evaluate the density of 1s in the incidence matrix in correspondence of which the transition from the sparse to the dense regime happens, let us consider the case $N\gg 1$: more formally, this amounts to consider the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$, i.e. $p=O(1/N)$. Since $p=T/NL$ and $h\equiv T/L$ remains finite, the probability for a generic hyperedge to be empty remains finite as well: consistently, the expected number of empty hyperedges becomes
$\lim_{N\to+\infty}\langle N_{0}\rangle=\lim_{N\to+\infty}L\left(1-\frac{h}{N}\right)^{N}=Le^{-h};$ (13)
upon imposing $Le^{-h}\leq 1$, i.e. that the expected number of empty hyperedges is at most 1, one derives what may be called filling threshold, corresponding to $h_{f}^{\text{RHM}}=\ln L$. In words, a value $p>p_{f}^{\text{RHM}}\equiv h_{f}^{\text{RHM}}/N=\ln L/N$ ensures that the expected number of empty hyperedges in our random hypergraph is strictly less than one. The calculation above complements the result derived in Barthelemy (2022) about the percolation threshold for hypergraphs (see also Appendix A).
III.1.2 2. Estimation of the number of parallel hyperedges
Let us, now, move to considering the issue of parallel hyperedges. By definition, two, parallel hyperedges $\alpha$ and $\beta$ are characterised by identical columns: hence, their Hamming distance, defined as the number of positions at which the corresponding symbols are different, is zero. More formally,
$d_{\alpha\beta}\equiv\sum_{i=1}^{N}[I_{i\alpha}(1-I_{i\beta})+I_{i\beta}(1-I_{i\alpha})],$ (14)
a sum whose generic addendum is 1 in just two cases: either $I_{i\alpha}=1$ and $I_{i\beta}=0$ or $I_{i\alpha}=0$ and $I_{i\beta}=1$. Since $d_{\alpha\beta}\sim\text{Bin}(N,2p(1-p))$, one finds that $P(d_{\alpha\beta}=0)=[1-2p(1-p)]^{N}$ and $\langle d_{\alpha\beta}\rangle=2p(1-p)N$, $\forall\>\alpha\neq\beta$. Notice that
$\lim_{T\to NL}P(d_{\alpha\beta}=0)$ $=\lim_{T\to 0}P(d_{\alpha\beta}=0)=1$ (15)
and
$\lim_{T\to NL}\langle d_{\alpha\beta}\rangle$ $=\lim_{T\to 0}\langle d_{\alpha\beta}\rangle=0$ (16)
as $(2T/NL)(1-T/NL)\to 0$ in both regimes: in words, the probability of observing parallel hyperedges progressively rises both as a consequence of having many 1s and as a consequence of having few 1s. Analogously for the expected Hamming distance.
Let us, now, evaluate the expected Hamming distance between any, two hyperedges $\alpha$ and $\beta$, within the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$, i.e. $p=O(1/N)$. Since $p=T/NL$ and $h\equiv T/L$ remains finite, one finds that
$\lim_{N\to+\infty}\langle d_{\alpha\beta}\rangle=\lim_{N\to+\infty}\frac{2h}{N}\left(1-\frac{h}{N}\right)N=2h;$ (17)
upon imposing $2h\geq 1$, i.e. that the expected Hamming distance between any, two hyperedges $\alpha$ and $\beta$ is at least 1, one derives what may be called resolution threshold, corresponding to $h_{r}^{\text{RHM}}=1/2$. In words, a value $p>p_{r}^{\text{RHM}}\equiv h_{r}^{\text{RHM}}/N=1/2N$ ensures that, on average, any, two hyperedges $\alpha$ and $\beta$ differ by at least one element (see also Appendix A).
III.1.3 3. Estimation of the overlap between hyperedges
Finally, let us move to considering the issue of overlapping hyperedges. To this aim, let us require that the probability for any, two hyperedges to overlap amounts to 1. In formulas,
$1-(1-p^{2})^{N}=1,$ (18)
a relationship leading to the equation
$(1-p^{2})^{N}\simeq 1-Np^{2}=0$ (19)
that holds true in the sparse regime and further implies
$h^{2}=N;$ (20)
such a value, firstly derived in Barthelemy (2022), individuates the so-called percolation threshold, corresponding to $h_{p}^{\text{RHM}}=\sqrt{N}$. In words, a value $p>p_{p}^{\text{RHM}}\equiv h_{p}^{\text{RHM}}/N=1/\sqrt{N}$ ensures that any, two hyperedges in our random hypergraph share at least one node.
Since $h_{r}^{\text{RHM}}\leq h_{f}^{\text{RHM}}\leq h_{p}^{\text{RHM}}$, such a result contributes to describe the following situation: by progressively rising the parameter $p$, three, different thresholds are met, i.e. i) the resolution threshold $p_{r}^{\text{RHM}}=1/2N$, ensuring that, on average, any, two hyperedges differ by at least one element; ii) the filling threshold $p_{f}^{\text{RHM}}=\ln L/N$, ensuring that the expected number of empty hyperedges is strictly less than one; iii) the percolation threshold $p_{p}^{\text{RHM}}=1/\sqrt{N}$, ensuring that any, two hyperedges overlap.
III.2 B. Heterogeneous benchmarks:
the Hypergraph Configuration Model (HCM)
The number of constraints can be enlarged to include the degrees, i.e. the sequence $\{k_{i}\}_{i=1}^{N}$, and the hyperdegrees, i.e. the sequence $\{h_{\alpha}\}_{\alpha=1}^{L}$. Counting the number of configurations on which both sequences match their empirical values is a hard task - although numerical recipes to shuffle the entries of a rectangular matrix, while preserving its marginals, exist Strona et al. (2014); Carstens (2015); Chodrow (2020). Solving the corresponding problem in the canonical framework is, instead, straightforward. Indeed, Shannon entropy maximisation leads to
$P(\mathbf{I})$ $=\frac{e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}$
$=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}$
$=\prod_{i=1}^{N}x_{i}^{k_{i}(\mathbf{I})}\prod_{\alpha=1}^{L}y_{\alpha}^{h_{\alpha}(\mathbf{I})}\prod_{i=1}^{N}\prod_{\alpha=1}^{L}(1+x_{i}y_{\alpha})^{-1},$ (21)
an expression that can be re-written as
$P(\mathbf{I})=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}p_{i\alpha}^{I_{i\alpha}}(1-p_{i\alpha})^{1-I_{i\alpha}}$ (22)
with $e^{-\alpha_{i}}\equiv x_{i}$, $\forall\>i$, $e^{-\beta_{\alpha}}\equiv y_{\alpha}$, $\forall\>\alpha$ and $p_{i\alpha}\equiv x_{i}y_{\alpha}/(1+x_{i}y_{\alpha})$, $\forall\>i,\alpha$. According to such a model, the entries of the incidence matrix of a hypergraph are independent random variables that obey different Bernoulli distributions, i.e. $I_{i\alpha}\sim\text{Ber}(p_{i\alpha})$, $\forall\>i,\alpha$. As a consequence, both degrees and hyperdegrees obey Poisson-Binomial distributions, i.e. $k_{i}\sim\text{PoissBin}(L,\{p_{i\alpha}\}_{\alpha=1}^{L})$, $\forall\>i$ and $h_{\alpha}\sim\text{PoissBin}(N,\{p_{i\alpha}\}_{i=1}^{N})$, $\forall\>\alpha$ Saracco et al. (2017).
In this case, solving the likelihood maximisation problem amounts to solve the system of coupled equations
$k_{i}^{*}$ $=\sum_{\alpha=1}^{L}\frac{x_{i}y_{\alpha}}{1+x_{i}y_{\alpha}}=\sum_{\alpha=1}^{L}p_{i\alpha}=\langle k_{i}\rangle,\>\forall\>i$ (23)
$h_{\alpha}^{*}$ $=\sum_{i=1}^{N}\frac{x_{i}y_{\alpha}}{1+x_{i}y_{\alpha}}=\sum_{i=1}^{N}p_{i\alpha}=\langle h_{\alpha}\rangle,\>\forall\>\alpha$ (24)
ensuring that $\langle k_{i}\rangle=k_{i}^{*}$, $\forall\>i$, $\langle h_{\alpha}\rangle=h_{\alpha}^{*}$, $\forall\>\alpha$ (and, as a consequence, $\langle T\rangle=T^{*}$). In case hypergraphs are sparse and in absence of hubs
$p_{i\alpha}\simeq x_{i}y_{\alpha}=\frac{k_{i}^{*}h_{\alpha}^{*}}{T^{*}},\>\forall\>i,\alpha.$ (25)
The HCM reduces to a ‘partial’ Configuration Model Saracco et al. (2017) when either the degree or the hyperdegree sequence is left unconstrained (see also Appendix B). The canonical ensemble of each randomisation model (Table 2 in Appendix B sums up the set of constraints defining them) can be explicitly sampled by considering each entry of $\mathbf{I}$, drawing a real number $u_{i\alpha}\in U[0,1]$ and posing $I_{i\alpha}=1$ if $u_{i\alpha}\leq p_{i\alpha}$, $\forall\>i,\alpha$.
The HCM is formally equivalent to the Bipartite Configuration Model (BiCM) Saracco et al. (2015). Such an identification is guaranteed by our focus on non-simple hypergraphs.
III.2.1 1. Estimation of the number of empty hyperedges
Let us, now, consider the probability for the generic hyperedge $\alpha$ to be empty or, in other terms, that its hyperdegree $h_{\alpha}$ is zero. Upon remembering that $h_{\alpha}\sim\text{PoissBin}(N,\{p_{i\alpha}\}_{i=1}^{N})$, one finds
$p_{0}^{\alpha}\equiv\prod_{i=1}^{N}(1-p_{i\alpha}).$ (26)
Such a quantity leads to the probability of observing at least one, empty hyperedge, now reading
$p_{1}\equiv 1-\prod_{\alpha=1}^{L}\left[1-p_{0}^{\alpha}\right]=1-\prod_{\alpha=1}^{L}\left[1-\prod_{i=1}^{N}(1-p_{i\alpha})\right],$ (27)
and to the expected number of empty hyperedges
$\langle N_{0}\rangle\equiv\sum_{\alpha=1}^{L}p_{0}^{\alpha}=\sum_{\alpha=1}^{L}\prod_{i=1}^{N}(1-p_{i\alpha}).$ (28)
As in the previous section, let us, now, inspect the behaviour of the aforementioned quantities on the ensemble induced by the HCM, as the density of 1s in the incidence matrix varies. Although it depends on (the heterogeneity of) the sets of coefficients $\{x_{i}\}_{i=1}^{N}$ and $\{y_{\alpha}\}_{\alpha=1}^{L}$, general conclusions can still be drawn within a simpler framework. To this aim, let us consider the functional form reading
$p_{i\alpha}=\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}},\>\forall\>i,\alpha$ (29)
where the vector of fitnesses $\{f_{i}\}_{i=1}^{N}$ accounts for the heterogeneity of nodes, the vector of fitnesses $\{g_{\alpha}\}_{\alpha=1}^{L}$ accounts for the heterogeneity of hyperedges and $z$ tunes the density of 1s in the incidence matrix [^1] Partial Configuration Models are recovered upon posing either $f_{i}=1$, $\forall\>i$ or $g_{\alpha}=1$, $\forall\>\alpha$.. Within such a framework, the fitnesses of the nodes and the fitnesses of the hyperedges can be drawn from any distribution. The dense and sparse regimes are, now, defined by the positions $z\to+\infty$ and $z\to 0$, respectively. In the dense case, one finds
$\lim_{z\to+\infty}p_{0}^{\alpha}$ $=\lim_{z\to+\infty}\prod_{i=1}^{N}(1-p_{i\alpha})$
$=\lim_{z\to+\infty}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=0,$ (30)
a relationship inducing $p_{1}\xrightarrow[]{z\to+\infty}0$ and $\langle N_{0}\rangle\xrightarrow[]{z\to+\infty}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of 1s increases. Consistently, in the sparse case one finds
$\lim_{z\to 0}p_{0}^{\alpha}$ $=\lim_{z\to 0}\prod_{i=1}^{N}(1-p_{i\alpha})$
$=\lim_{z\to 0}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=1,$ (31)
a relationship inducing $p_{1}\xrightarrow[]{z\to 0}1$ and $\langle N_{0}\rangle\xrightarrow[]{z\to 0}L$. In words, the probability of observing empty hyperedges progressively rises as the density of 1s decreases.
III.2.2 2. Estimation of the number of parallel hyperedges
Let us, now, focus on the issue of parallel hyperedges. As in the case of the RHM, we consider the Hamming distance between the columns representing the two hyperedges $\alpha$ and $\beta$. Since, now, $d_{\alpha\beta}\sim\text{PoissBin}(N,\{q_{i}^{\alpha\beta}\}_{i=1}^{N})$, where $q_{i}^{\alpha\beta}\equiv p_{i\alpha}(1-p_{i\beta})+p_{i\beta}(1-p_{i\alpha})$ with $p_{i\alpha}=zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$ and $p_{i\beta}=zf_{i}g_{\beta}/(1+zf_{i}g_{\beta})$, one finds that $P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$ and $\langle d_{\alpha\beta}\rangle=\sum_{i=1}^{N}q_{i}^{\alpha\beta}$, $\forall\>\alpha\neq\beta$. Notice that
$\lim_{z\to+\infty}P(d_{\alpha\beta}=0)$ $=\lim_{z\to 0}P(d_{\alpha\beta}=0)=1$ (32)
and
$\lim_{z\to 0}\langle d_{\alpha\beta}\rangle$ $=\lim_{z\to 0}\langle d_{\alpha\beta}\rangle=0$ (33)
as $q_{i}^{\alpha\beta}\to 0$ in both regimes. As in the case of the RHM, the probability of observing parallel hyperedges progressively rises both as a consequence of having many 1s and as a consequence of having few 1s. Analogously for the expected Hamming distance.
When considering heterogeneous benchmarks, closed-form expressions for the asymptotic estimates can be provided only in specific cases (see also Appendix B).
Figure 2: Left panel: trend of $p_{1}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$, i.e. the probability of observing at least one empty hyperedge as a function of $T/NL$; evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ (whose related, critical density value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$ individuates a steep transition separating the sparse regime from the dense one) returns the value $0.63231-(1-1/L)^{L}=$ 0.6323. Right panel: trend of $p_{1}$ as defined by Eq. (27). Notice that the trend induced by the HCM is shifted on the right with respect to the trend induced by the RHM: this is probably due to the presence of small fitnesses increasing the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $p_{1}$ vanish. The dense (sparse) regime is recovered for large (small) values of $T/NL$. Each dot represents an average taken over an ensemble of $1000$ configurations (explicitly sampled from either the RHM or the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
IV IV. Results
IV.1 A. Hypergraphs in the dense and sparse regime
Let us start by verifying the correctness of the estimations of the number of empty and parallel hyperedges provided by our benchmarks: to this aim, we have considered the values $N=300$ and $L=1000$ (for other values, see Appendix A and Appendix B).
IV.1.1 1. The Random Hypergraph Model
Let us start focusing on our homogeneous benchmark. The left panel of Fig. 2 depicts the trend of $p_{1}=1-\left[1-\left(1-\frac{T}{NL}\right)^{N}\right]^{L}$. The left panel of Fig. 3 depicts the trend of $\frac{\langle N_{0}\rangle}{L}=\left(1-\frac{T}{NL}\right)^{N}=p_{0}$ and $P(d_{\alpha\beta}=0)=\left[1-\frac{2T}{NL}\left(1-\frac{T}{NL}\right)\right]^{N}$. Each quantity has been plotted as a function of $T/NL\in[10^{-6},1]$. The dense (sparse) regime is recovered for large (small) values of $T/NL$. Each dot of Figs. 2 and 3 represents an average taken over an ensemble of $1000$ configurations explicitly sampled from the RHM and is accompanied by the corresponding $95\%$ confidence interval. A first, general comment concerns the agreement between the analytical trends (solid lines) and the numerical estimations (dots), confirming the correctness of our formulas.
The filling threshold.
Although the value of the filling threshold has been determined by inspecting the asymptotic behaviour of $\langle N_{0}\rangle$, the quantity showing the neatest transition from the sparse to the dense regime is $p_{1}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$; evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $1-(1-1/L)^{L}\simeq 0.6323$ interestingly enough, letting $L$ grow asymptotically as well, the previous value tends to $1-e^{-1}\simeq 0.6321$.
Upon considering that $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $0.0011/L=$ 0.001: in words, although the filling threshold ensures that each, single hyperedge is empty with an overall small probability, the likelihood of observing at least one, empty hyperedge is still large (i.e. 2/3$\simeq$ 2/3)≃ 2 / 3 ). The steepness of the trend of $p_{1}$, however, suggests it to quickly vanish as the density of 1s in the incidence matrix crosses the value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$.
Figure 3: Left panel: trends of $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$ (aquamarine line) and $P(d_{\alpha\beta}=0)\xrightarrow[]{N\to+\infty}e^{-2h}$ (dark-blue line) as a function of $T/NL$; evaluating them in correspondence of $h_{r}^{\text{RHM}}=1/2$ (the related, critical density value is indicated in light-blu) returns the values $e^{-{h_{r}^{\text{RHM}}}}=1/\sqrt{e}\simeq 0.606$ and $e^{-{2h_{r}^{\text{RHM}}}}=e^{-1}\simeq 0.367$; evaluating them in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ (the related, critical density value is indicated in magenta) returns the values $e^{-{h_{f}^{\text{RHM}}}}=1/L=0.001$ and $e^{-{2h_{f}^{\text{RHM}}}}=1/L^{2}=10^{-6}$. The red line indicates the density value corresponding to $h=1$, ensuring that each, expected hyperdegree amounts to 1. Right panel: trends of $p_{0}$, as defined by Eq. (26), and $P_{0}$, i.e. the percentage of pairs of parallel hyperedges. As already noticed, the shift on the right of the trend induced by the HCM is probably due to the presence of small fitnesses, decreasing the magnitude of the set of probability coefficients. The dense (sparse) regime is recovered for large (small) values of $T/NL$. Each dot represents an average taken over an ensemble of $1000$ configurations (explicitly sampled from either the RHM or the HCM) and is accompanied by the corresponding $95\%$ confidence interval. The yellow line indicates the density value in correspondence of which the minimum, expected hyperdegree amounts to 1.
The resolution threshold.
The value of the resolution threshold has been determined by inspecting the asymptotic behaviour of $\langle d_{\alpha\beta}\rangle$: evaluating $P(d_{\alpha\beta}=0)\xrightarrow[]{N\to+\infty}e^{-2h}$ in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $e^{-1}\simeq 0.367$.
Upon considering that $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating it in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $1/\sqrt{e}\simeq 0.606$: in words, the resolution threshold ensures the likelihood of observing any, two parallel hyperedges to be almost half of the probability of each, single hyperedge to be empty, hence suggesting that getting rid of parallel hyperedges is ‘easier’ than getting rid of empty ones.
Crossing the thresholds.
As Fig. 3 confirms, the probability for any, two hyperedges to be parallel decreases faster than the probability of each, single hyperedge to be empty: in words, while filling a large portion of the incidence matrix is not required to observe a limited amount of parallel hyperedges, when empty hyperedges are no longer observed, parallel hyperedges are no longer observed, too. More formally, $h_{r}^{\text{RHM}}=1/2<h_{f}^{\text{RHM}}=\ln L$: hence, progressively rising the density of 1s in the incidence matrix leads us crossing the resolution threshold before crossing the filling threshold. In more quantitative terms, evaluating $\langle N_{0}\rangle\xrightarrow[]{N\to+\infty}Le^{-h}$ in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $L/\sqrt{e}\simeq 606$ which is larger than $\langle d_{\alpha\beta}\rangle\xrightarrow[]{N\to+\infty}2h$ evaluated in correspondence of $h_{r}^{\text{RHM}}=1/2$ returning the value 1 by definition.
IV.1.2 2. The Hypergraph Configuration Model
Let us, now, focus on the trends induced by our heterogeneous benchmarks. Following the procedure described in the previous sections, both the fitnesses of nodes and those of hyperedges were drawn from a Pareto distribution [^2] Other fat-tailed distributions were considered: qualitatively, results do not change. with $\alpha=2$. The right panel of Fig. 2 depicts the trend of $p_{1}$, as defined by Eq. (27). The right panel of Fig. 3 depicts the trend of $p_{0}$, as defined by Eq. (26), and the arithmetic mean $P_{0}\equiv 2\sum_{\alpha<\beta}P(d_{\alpha\beta}=0)/L(L-1)$ with $P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$, i.e. the average of the value $P(d_{\alpha\beta}=0)$ over all pairs of hyperedges - a choice dictated by the will of accounting for the heterogeneity of different pairs of hyperedges. Both quantities have been plotted as a function of $\langle T\rangle/NL\in[10^{-6},1]$ where $\langle T\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$. The dense (sparse) regime is recovered for large (small) values of $z$. Each dot of Figs. 2 and 3 represents an average taken over an ensemble of $1000$ configurations explicitly sampled from the HCM and is accompanied by the corresponding $95\%$ confidence interval. As in the case of the RHM, the agreement between the analytical trends (solid lines) and the numerical estimations (dots) confirms the correctness of our formulas.
Comparing the analytical estimation of (the trend of) the percentage of pairs of parallel hyperedges with its sample counterpart reveals the former to provide a quite accurate estimation of the latter. Besides, we explicitly notice that the trends characterising the HCM are sistematically shifted on the right: this is probably due to the presence of small fitnesses that, at least in case of Fig. 2, increase the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $p_{1}$ vanish. Such a result seems to be confirmed by the trends depicted in Fig. 3.
Deriving an explicit expression for the filling threshold in the case of the HCM is a rather difficult task; still, we can proceed in a purely numerical fashion and individuate the value of the density of 1s in the incidence matrix guaranteeing that the minimum, expected hyperdegree amounts to 1: as evident from Fig. 3, such a threshold still lies in the right tail of the trends induced by the HCM - although its, precise, numerical value depends on those of the fitnesses.
IV.2 B. Solving the HCM on real-world hypergraphs
In order to test our benchmarks on real-world configurations, we have focused on a number of data sets taken from Austin R. Benson’s website [^3] https://www.cs.cornell.edu/~arb/data/, i.e. the contact-primary-school, the email-Enron and the NDC-classes ones.
To have an explicit expression for the expected value of any quantity of interest, let us comment on the goodness of the approximation
$p_{i\alpha}\simeq x_{i}y_{\alpha}=\frac{k_{i}^{*}h_{\alpha}^{*}}{T^{*}},\>\forall\>i,\alpha.$ (34)
As Fig. 7 in Appendix A shows, it is is quite accurate for each data set considered here - in fact, one can safely assume that $x_{i}\simeq k_{i}^{*}/\sqrt{T^{*}}$, $\forall\>i$ and $h_{\alpha}^{*}/\sqrt{T^{*}}$, $\forall\>\alpha$.
IV.3 C. ‘Hypergraph to graph’ projection
The canonical formalism that we have adopted leads to factorisable distributions, i.e. distributions that can be written as a product of pair-wise probability distributions; this allows the expectation of several quantities of interest to be evaluated analytically. To this aim, let us consider the matrix introduced in Battiston et al. (2020), reading
$\mathbf{W}=\mathbf{I}\cdot\mathbf{I}^{T}-\mathbf{K}$ (35)
with $\mathbf{K}$ being the diagonal matrix whose $i$-th entry reads $k_{i}$; according to the definition above
$w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}-\delta_{ij}k_{i}$ (36)
counts the number of hyperedges nodes $i$ and $j$ are connected by - more explicitly, $w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$, $i\neq j$ and $w_{ii}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{i\alpha}-k_{i}=\sum_{\alpha=1}^{L}I_{i\alpha}-k_{i}=k_{i}-k_{i}=0$. The null models discussed so far can be employed to calculate $\langle w_{ij}\rangle$, $i\neq j$ that, in a perfectly general fashion, reads
$\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}\langle I_{i\alpha}I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}\langle I_{i\alpha}\rangle\langle I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}p_{j\alpha};$ (37)
the total number of hyperedges shared by node $i$ with any other node in the hypergraph (in a sense, its ‘strength’ - see also Fig. 4) can be computed as
$\sigma_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}w_{ij}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$ (38)
whose expected value reads
$\langle\sigma_{i}\rangle$ $=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}[\langle h_{\alpha}\rangle-p_{i\alpha}].$ (39)
As further confirmed by Fig. 8 in Appendix A, the approximation provided by Eq. (34) allows us to pose $\langle\sigma_{i}\rangle_{\text{HCM}}\simeq k_{i}^{*}\sum_{\alpha=1}^{L}(h_{\alpha}^{*})^{2}/T^{*}$, $\forall\>i$. Interestingly, as Fig. 5a shows, the HCM overestimates the extent to which any two nodes of the email-Enron data set overlap: in words, such a real-world hypergraph is more compartmentalised than expected.
Let us, now, extend the concept of assortativity to hypergraphs. To this aim, we consider the quantity named average incident hyperedges degree, defined as
$k_{i}^{nn}=\sum_{\alpha=1}^{L}\frac{I_{i\alpha}h_{\alpha}}{k_{i}}=\frac{\sigma_{i}+k_{i}}{k_{i}}=\frac{\sigma_{i}}{k_{i}}+1\simeq\frac{\sigma_{i}}{k_{i}}$ (40)
and representing the arithmetic mean of the degrees of the hyperedges including node $i$. An analytical approximation of its expected value can be provided as well:
$\langle k_{i}^{nn}\rangle$ $\simeq\sum_{\alpha=1}^{L}\frac{p_{i\alpha}[\langle h_{\alpha}\rangle+1-p_{i\alpha}]}{\langle k_{i}\rangle}$
$=\frac{\langle\sigma_{i}\rangle+\langle k_{i}\rangle}{\langle k_{i}\rangle}=\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}+1\simeq\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}.$ (41)
Figure 4: Graphical representation of the ‘hypergraph to graph’ projection. Notice that while the degree of each node (i.e. the number of hyperedges that are incident to it) reads $\{k_{i}\}=(3,3,3,3,4,1)$, the sigma of each node (i.e. the total number of hyperedges shared by it with any other node or, equivalently, the strength on the projection) reads $\{\sigma_{i}\}=(8,8,9,6,10,3)$ and the kappa of each node (i.e. the number of nodes it shares an hyperedge with or, equivalently, the degree on the projection) reads $\{\kappa_{i}\}=(5,4,5,4,5,3)$.
IV.4 D. Disparity ratio and degree in the projection
More information about the patterns shaping real-world hypergraphs can be obtained upon defining the ratio $f_{ij}=w_{ij}/\sigma_{i}$, $i\neq j$ that induces the quantity
$Y_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}f_{ij}^{2}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\frac{w_{ij}^{2}}{\sigma_{i}^{2}},$ (42)
known as disparity ratio and quantifying the (un)evenness of the distribution of the weights constituting the strength of node $i$ over the $\kappa_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\Theta[w_{ij}]\equiv\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}a_{ij}$ links characterising its connectivity - since $a_{ij}=1$ if nodes $i$ and $j$ share, at least, one hyperedge, $\kappa_{i}$ is the degree of node $i$ in the projection of the hypergraph (see also Fig. 4 and Fig. 5b). Since, under the RHM, $w_{ij}\sim\text{Bin}(L,p^{2})$, we find that $\langle a_{ij}\rangle=1-(1-p^{2})^{L}$, i.e. the expected value of $a_{ij}$ coincides with the probability of observing a non-zero overlap. Under the HCM, instead, $w_{ij}\sim\text{PoissBin}(L,\{p_{i\alpha}p_{j\alpha}\}_{\alpha=1}^{L})$, hence
$\langle a_{ij}\rangle=1-\prod_{\alpha=1}^{L}(1-p_{i\alpha}p_{j\alpha}).$ (43)
Let us also notice that
$Y_{i}=\frac{1}{\kappa_{i}}$ (44)
in case weights are equally distributed among the connections established by node $i$, i.e. $w_{ij}=a_{ij}\sigma_{i}/\kappa_{i}$, $i\neq j$. Any larger value signals an excess concentration of weight in one or more links. An analytical approximation of the expected value of the disparity ratio of node $i$ can be provided as well:
$\langle Y_{i}\rangle\simeq\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\frac{\langle w_{ij}^{2}\rangle}{\langle\sigma_{i}^{2}\rangle}.$ (45)
Figure 5: Scatter plots between the empirical values $\{\sigma_{i}\}$, $\{\kappa_{i}\}$, $\{Y_{i}\}$, $\{\text{CEC}_{i}\}$ and the expected values $\{\langle\sigma_{i}\rangle_{\text{HCM}}\}$, $\{\langle\kappa_{i}\rangle_{\text{HCM}}\}$, $\langle\{Y_{i}\rangle_{\text{HCM}}\}$, $\{\langle\text{CEC}_{i}\rangle_{\text{HCM}}\}$. The HCM overestimates the extent to which any, two nodes overlap, as well as the CEC; the disparity ratio, instead, is underestimated by it. These results can be understood by considering that the HCM just constrains the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed.
Contrarily to what has been previously observed, the expected value of the disparity ratio cannot be always safely decomposed as a ratio of expected values - not even if the ‘full’ HCM is employed. In fact, while this approximation works relatively well for the contact-primary-school data set, it does not for the email-Enron and the NDC-classes ones (see also Fig. 10 in Appendix C). For this reason, the expected value of the disparity ratio has been evaluated by explicitly sampling the ensemble of incidence matrices induced by the ‘full’ HCM. In any case, as Fig. 5c shows, such a null model underestimates the disparity ratio characterising each node of the email-Enron data set: in words, the empirical overlap between any two nodes is (much) less evenly ‘distributed’ than expected. A similar conclusion can be drawn by considering $\langle\kappa_{i}\rangle=\sum_{i=1}^{N}[1-\prod_{\alpha=1}^{L}(1-p_{i\alpha}p_{j\alpha})]$: as Fig. 5b shows, the degree of the nodes in the projection tend to be significantly smaller than expected, meaning that hyperedges concentrate on fewer edges than expected. This observation is in line with recent works showing the encapsulation and ‘simpliciality’ of real-world hypergraphs LaRock and Lambiotte (2023); Landry et al. (2024).
IV.5 F. Eigenvector centrality
Centrality measures for hypergraphs have been defined as well. An example is provided by the clique motif eigenvector centrality (CEC), defined in Benson (2019) (see also Appendix D): $\text{CEC}_{i}$ corresponds to the $i$-th entry of the Perron-Frobenius eigenvector of $\mathbf{W}$. As Fig. 5d shows, the HCM underestimates the CEC as well: such a result can be understood by considering that the HCM constrains only the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed, an evidence letting the nodes overlap more, thus causing the entries of $\langle\mathbf{W}\rangle$ to be overall larger and less dissimilar, as well as those of its Perron-Frobenius eigenvector.
Figure 6: Top panels: projections of the contact-primary-school, email-Enron and NDC-classes data sets onto the layer of nodes. Bottom panels: validated counterparts of the aforementioned projections: any two nodes are linked if they share a significantly large number of hyperedges. Communities have been detected by running the Louvain algorithm.
IV.6 G. Confusion matrix
Let us, now, consider the set of properties composing the so-called confusion matrix (see also Appendix E). They are known as the true positive rate (TPR), i.e. the percentage of 1s correctly recovered by a given method and whose expected value reads
$\langle\text{TPR}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{T};$ (46)
the specificity (SPC), i.e. the percentage of 0s correctly recovered by a given method and whose expected value reads
$\langle\text{SPC}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{(1-I_{i\alpha})(1-p_{i\alpha})}{NL-T};$ (47)
the positive predictive value (PPV), i.e. the percentage of 1s correctly recovered by a given method with respect to the total number of 1s predicted by it and whose expected value reads
$\langle\text{PPV}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{\langle T\rangle};$ (48)
and the accuracy (ACC), measuring the overall performance of a given method in correctly placing both 1s and 0s and whose expected value reads
$\langle\text{ACC}\rangle=\frac{\langle\text{TP}\rangle+\langle\text{TN}\rangle}{NL}.$ (49)
Results on the confusion matrix of a number of real-world hypergraphs reveal that the large sparsity of the latter ones makes it difficult to reproduce the TPR and the PPV (see also Table 3 in Appendix E); on the other hand, the capability of the HCM (both in its ‘full’ and approximated version) to reproduce the density of 1s - and, as a consequence, the density of 0s - ensures the SPC to be recovered quite precisely, in turn ensuring the overall accuracy of the model to be large (for an overall evaluation of the performance of the HCM in reproducing real-world hypergraphs, see also Table 4 in Appendix F).
IV.7 H. Community detection
Communities are commonly understood as densely connected groups of nodes. Representing an hypergraph via its incidence matrix allows this statement to be made more precise from a statistical perspective: in fact, the null models discussed so far can be employed to test if any two nodes share a significantly large number of hyperedges - hence can be clustered together, should this be the case.
To this aim, we can adapt the recipe proposed in Saracco et al. (2017) to project bipartite networks and summarised in the following. One, first, computes
[…]
where $n=N(N-1)/2$ and $t$ is the single-test significance level, set to $0.01$ in the present analysis. Third, one links the (pairs of) nodes whose related p-value is smaller than the aforementioned threshold.
Figure 6 shows the result returned by the Louvain algorithm, run on the validated projections: as noticed elsewhere Saracco et al. (2017); Pratelli et al. (2024); Guarino et al. (2024), the detection of mesoscale structures is enhanced if carried out on filtered topologies.
V V. Conclusions
Our paper contributes to current research on hypergraphs by extending the constrained entropy-maximisation framework to incidence matrices, i.e. their simplest, tabular representation. Differently from the currently-available techniques Chodrow (2020), our methodology has the advantage of being analytically tractable, scalable and versatile enough to be straightforwardly extensible to directed and/or weighted hypergraphs.
Beside leading to results whose relevance is mostly theoretical (i.e. the individuation of different regimes for higher-order structures and the estimation of the actual impact of empty and parallel hyperedges on the analysis of empirical systems), our models prove to be particularly useful when employed as benchmarks for real-world systems, i.e. for detecting patterns that are not imputable to purely random effects. Specifically, our results suggest that real-world hypergraphs are characterised by a degree of self-organisation that is absolutely non-trivial (see also Appendix F).
This is even more surprising when considering that our results are obtained under a benchmark such as the HCM, i.e. a null model constraining both the degree and the hyperdegree sequences: since it overestimates the extent to which any two nodes overlap - a result whose relevance becomes evident as soon as one considers the effects that higher-order structures have on spreading and cooperation processes de Arruda et al. (2021); Iacopini et al. (2022); St-Onge et al. (2022) - our future efforts will be directed towards the analysis of benchmarks constraining non-linear quantities such as the co-occurrences between nodes and/or hyperedges.
VI VI. Acknowledgements
RL acknowledges support from the EPSRC grants EP/V013068/1, EP/V03474X/1 and EP/Y028872/1. TS acknowledges support from SoBigData.it that receives funding from European Union – NextGenerationEU – National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) – Project: ‘SoBigData.it – Strengthening the Italian RI for Social Mining and Big Data Analytics’ - Prot. IR0000013 – Avviso n. 3264 del 28/12/2021. FS acknowledges support from the project ‘CODE - Coupling Opinion Dynamics with Epidemics’, funded under PNRR Mission 4 ‘Education and Research’ - Component C2 - Investment 1.1 - Next Generation EU ‘Fund for National Research Program and Projects of Significant National Interest’ PRIN 2022 PNRR, grant code P2022AKRZ9.